A chemist needs 180 milliliters of a 39% solution but has only 15% and 69% solutions available. How many milliliters of each should be mixed to obtain the desired solution?
Answer by TimothyLamb(4379)

x = ml of 15%
y = ml of 69%
---
15x + 69y = 39*180
x + y = 180
---
put the system of linear equations into standard form
---
15x + 69y = 7020
x + y = 180
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = ml of 15% = 100
y = ml of 69% = 80
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php